In testing the strength of chains under heat treatments A and B, the following two sets of strengths were found.

A : 3610 , 3608 , 3672

B : 3416 , 3460 , 3390

a) Find a 90 % confidence interval for \sigma_a ^2 /\sigma_b ^2 . State your assumption(s) for your answers to be valid.

b) On the basis of the confidence interval in part (a), do you think it is likely that

\sigma_a ^2 =\sigma_b ^2 ?Why?
Answer #1

For A, sa2 = 1324

For B, sb2 = 1252

a) At 90% confidence intrval the critical values are F(\small \alpha/2, n_{a} - 1, n_{b} - 1) = 19

                                                                                 F(\small 1 - \alpha/2, n_{a} - 1, n_{b} - 1) = 0.0526

The 90% confidence interval for \small \sigma_{a}^2/{\sigma_{b}^2} is

(sa2/sb2) * (1/F(\small \alpha/2, n_{a} - 1, n_{b} - 1)) < \small \sigma_{a}^2/{\sigma_{b}^2} < (sa2/sb2) * (1/F(\small 1 - \alpha/2, n_{a} - 1, n_{b} - 1)

= (1324/1252) * (1/19) < \small \sigma_{a}^2/{\sigma_{b}^2} < (1324/1252) * (1/0.0526 )

=   0.0557 < \small \sigma_{a}^2/{\sigma_{b}^2} < 20.1047

b) Since the confidence interval cointans 1, so we can be conclude that \small \sigma_{a}^2 = {\sigma_{b}^2}